40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 32     Entry time: Tue Feb 26 19:52:08 2008
Author: Rana 
Type: Electronics 
Category: CDS 
Subject: channels 
Stefan, Rana

We changed the cable hookup -- to check out the AI box + ADC.
Ch1 = 3 Vpp offset from a SRS function generator    C2:OMS-SUS_TOP1_OUT_2048
Ch2 = 50 Ohms                                       C2:OMS-SUS_TOP2_OUT_2048
Ch6 = ISS SENSE PD DC output                        C2:OMS-SUS_SIDE_OUT_2048

The laser is now on and set to record the power fluctuations over a long term. The value of the SIDE input gain
is set such that SUS_SIDE_OUT16 is in units of mW on the Ophir high power head.

--------------

We also noticed that the signals coming out of the AI box are very aliased. The thing is running at 2048 Hz but we think
that the AA and AI filters have 10 kHz cutoffs (D070081-00).

Next:
-----
- Redo the FE diagram to get rid of OMC SUS and make a PSL diagram.
- Get the trending working right.
- Up the FE sample rate to 32768 Hz
ELOG V3.1.3-